# Inertial range of magnetorotational turbulence

### v1

@@ Discussion
The validity of reduced MHD in solving the inertial range of magnetorotational turbulence is now supported by the two findings of this study, namely (1) that the spatial structures of the small-scale fluctuations in our magnetorotational turbulence are elongated along the large-scale magnetic field which is azimuthally elongated and (2) that their spectral shapes remarkably resemble those obtained by reduced MHD. Although the results of our MHD simulation are formally applicable when $\beta$, thermal-to-magnetic pressure ratio, is infinitely large as we solved incompressible MHD, the simulations of reduced MHD showed that the spectral shape and the energy partition between the slow-magnetosonic and Alfvénic fluctuations (viz., by the factor of two) do not depend on $\beta$. Thus, we think that the results shown in this paper are also valid in the smaller $\beta$ regime. Note that the recent study shows that the ratio of energy injection between the slow-magnetosonic and Alfvénic fluctuations does not depend on $\beta$  (?), which supports our claim.
@@
The energy partition between the slow-magnetosonic and Alfvénic fluctuations is important for inferring the ion vs. electron heating in hot accretion flows, such as at M87 and Sgr A*. The ion-to-electron heating ratio is a key parameter for theoretical understanding of radio observations by the Event Horizon Telescope  (?). In collisionless magnetorotational turbulence, the half of the energy flux injected via MRI is supposed to be viscously dissipated due to the pressure anisotropy  (?, ?), and the remaining half turns into ion and electron heating at the microscopic scales smaller than the ion Larmor radius. Our previous study of microscopic turbulence using hybrid gyrokinetics showed that, regarding the heating at the microscopic scales, ions are heated more efficiently than electrons when slow-mode-like compressive fluctuations dominate the Alfvénic ones  (?). Thus, the fact that the slow-magnetosonic fluctuations have twice as large energy as the Alfvénic ones indicates that at least half of the energy cascaded down to the ion Lamor scale dissipates into ion heating (see also  (?) for some caveats of this conclusion). Recently, the one dimensional energy transport model of hot accretion disks showed that the preferential ion heating due to the slow-mode-like compressive fluctuations can significantly influence the global temperature distribution  (?). We should note that this discussion assumes that MRI turbulence is active in the accretion flows around M87 and Sgr A*  (?, ?). However, the theoretical examinations of the Event Horizon Telescope data prefer magnetically arrested disk (MAD) regime, in which MRI is suppressed due to strong magnetic fields.
@@
MRI turbulence can accelerate high-energy nonthermal protons, and these protons can produce cosmic neutrinos seen by IceCube experiment  (?). Previous studies on stochastic acceleration using MHD and test-particle simulations exhibit particle acceleration with a hard-sphere type diffusion coefficient in momentum space  (?, ?, ?). However, these MHD simulations lack sufficient spacial resolution, underestimating the acceleration efficiency at lower energies. Our highest resolution MRI simulation resolves the inertial range of MRI turbulence for the first time. This allows us to evaluate the particle acceleration efficiency at much lower energies down to dissipation scale. This would shed light on modeling particle acceleration inside accretion flows.
@@
Acknowledgments
YK thanks Hideaki Miura, Tomohiko Watanabe, and Jim Stone for fruitful comments on this study. YK is supported by JSPS KAKENHI grant No. 20K14509. SSK is supported by JSPS KAKENHI grant Nos. 22K14028, 21H04487, 23H04899 and the Tohoku Initiative for Fostering Global Researchers for Interdisciplinary Sciences (TI-FRIS) of MEXTs Strategic Professional Development Program for Young Researchers. Numerical computations reported here were carried out on Fugaku at the RIKEN Center for Computational Science (Project ID: hp220027 and hp230006), on ATERUI II at Centre for Computational Astrophysics in National Astronomical Observatory of Japan, on Oakbridge-CX, Oakforest-PACS and Wisteria/BDEC-01 at the University of Tokyo, on Flow at Nagoya University, and on ITO at Kyushu University.
Figure 1: Morphology of finely resolved spatial structures of magnetorotational turbulence. Spatial distributions of (A) flow and (B) magnetic field intensity on $x=0$, $y=0$, and $z=0$ planes where $x,y$, and $z$ denote radial, azimuthal, and vertical directions, respectively. (C) and (D) depict the flow and magnetic field intensity distributions, respectively, high-pass filtered at $|\mathbf{k}|\geq 20$ (as shown in Fig. 2C, the fluctuations with $|\mathbf{k}|\geq 20$ are in the inertial range). The bluish pseudo color in (C) indicates the magnitude of unfiltered electric current density $\mathbf{J}$ in a logarithmic scale. The green lines are typical magnetic field lines.
Figure 2: Spectra of magnetorotational turbulence. (A) Kinetic, magnetic, and total energy spectra compensated by $k^{3/2}$, where $k$ is the wavenumber. (B) Spectra of $x$, $y$, and $z$ components of magnetic energy. (C) Cross-scale energy flux $\Pi^{f^{<}}_{g^{>}}$ denoting the transfer from the field $f$ (which is the flow field when $f=u$ and is the magnetic field when $f=B$) with the wavenumber smaller than $k$ to the field $g$ with the wavenumber larger than $k$. The gray shaded area indicates to the dissipation range.
Figure 3: Locality of nonlinear energy transfer. (A)-(D) Shell-to-Shell energy transfer function $\mathcal{T}{fg}$, representing the transfer from field $f$ (the flow field for $f=u$ and the magnetic field for $f=B$) with wavenumber $Q$ to field $g$ with wavenumber $K$. (E)-(H) Contribution of the energy transfer from the injection range $k\leq 4$.
Supplementary Materials for
Inertial range of magnetorotational turbulence
Yohei Kawazura, Shigeo S. Kimura
*Correspondence to: kawazura@tohoku.ac.jp
Materials and Methods
@@ Acknowledgments
$\frac{\partial\mathbf{u}}{\partial t}+(\mathbf{u}_{0}+\mathbf{u})\cdot\nabla\mathbf{u}=-\nabla P+\mathbf{B}\cdot\nabla\mathbf{B}-2\Omega\hat{z}\times\mathbf{u}-\mathbf{u}\cdot\nabla\mathbf{u}_{0}$ (1)
@@
where $\mathbf{u}$ is the flow velocity, $\mathbf{B}$ is the magnetic field, $P$ is the total pressure, $\Omega$ is the local angular velocity of the disk, $q=-(\mathrm{d}\ln\Omega/\mathrm{d}\ln r)_{r=r_{0}}$ is the shear rate, and $\mathbf{u}_{0}=-q\Omega x\hat{\mathbf{y}}$ is the background flow. Due to the incompressible condition, the density $\rho$ is spatio-temporally constant. The incompressible condition is applicable to high beta (thermal-to-magnetic energy ratio) plasmas like those in accretion disks. In fact, the particle-in-cell simulation demonstrated that the incompressible approximation is valid in collisionless magnetorotational turbulence in a shearing box  (?). We assume that the rotation of the disk is Keplerian, i.e., $q=3/2$. The boundary condition is set to periodic in $y$ and $z$, and to shearing periodic in $x$  (?).
@@
We numerically solve (1)-(3) via pseudo-spectral code Calliope  (?). In order to adopt the pseudospectral method, we enforce the triply periodic boundary conditions by transforming to the shearing coordinate, $y\mapsto y-q\Omega tx$. This transformation makes the radial wavenumber time dependent, $k_{x}\mapsto k_{x}+q\Omega tk_{y}$. To avoid $k_{x}$ from ever growing, we adopt the remapping method where the fields are mapped to the original non-shearing coordinate every $T=L_{y}/(q\Omega L_{x})$  (?, ?).
@@
The computational domain was discretized into $N_{x}\times N_{y}\times N_{z}=N\times N\times N/2$ grid points. We set $N$ at 256 initially and gradually increased it to 8192 after nonlinear saturation. Each time $N$ was increased, we continued the simulation until the spectral shapes near the dissipation scale did not change before $N$ was increased again. When $N$ reached 8192, the simulation was continued for more than 100 $\Omega^{-1}$ from the beginning.
@@
To broaden the inertial range, we used fourth-order hyperviscosity $\nu_{\mathrm{h}}\nabla^{4}\mathbf{u}$ and hyperresistivity $\eta_{\mathrm{h}}\nabla^{4}\mathbf{B}$ to terminate the cascade. Magnetic Prandtl number $\mathrm{Pm}=\nu_{\mathrm{h}}/\eta_{\mathrm{h}}$ was set to unity. We found that when the order of the hyper-dissipation was eight, there appeared an unphysical roll up in the kinetic energy spectrum.
@@
Figure S1: Two dimensional slices of $x-$, $y-$, and $z-$components of $\mathbf{u}$ and $\mathbf{B}$.
Figure S2: Time evolution of spectra.

### v2

@@ Discussion
The validity of reduced MHD in solving the inertial range of magnetorotational turbulence is now supported by the two findings of this study, namely (1) that the spatial structures of the small-scale fluctuations in our magnetorotational turbulence are elongated along the large-scale magnetic field which is azimuthally elongated and (2) that their spectral shapes remarkably resemble those obtained by reduced MHD. Although the results of our MHD simulation are formally applicable when $\beta$, thermal-to-magnetic pressure ratio, is infinitely large as we solved incompressible MHD, the simulations of reduced MHD showed that the spectral shape and the energy partition between the slow-magnetosonic and Alfvénic fluctuations (viz., by the factor of two) do not depend on $\beta$. Thus, we think that the results shown in this paper are also valid in the smaller $\beta$ regime. Note that the recent study shows that the ratio of energy injection between the slow-magnetosonic and Alfvénic fluctuations does not depend on $\beta$  (?), which supports our claim. Apart from $\beta$, our results can depend on the value of Pm and on the presence or absence of net magnetic flux. While we have explored only the case with Pm = 1, with finite vertical net flux, and without azimuthal net flux, numerous simulations have investigated other cases (e.g., see Ref.  (?) for the latest study). However, these previous parameter scans mostly focus on the dependence of the value of $\alpha$ and saturation amplitude. Therefore, it would be very interesting to investigate whether the results of this study (e.g., the partition between slow-magnetosonic and Alfvénic fluctuations) change under different settings.
@@
The energy partition between the slow-magnetosonic and Alfvénic fluctuations is important for inferring the ion vs. electron heating in hot accretion flows, such as at M87 and Sgr A*. The ion-to-electron heating ratio is a key parameter for theoretical understanding of radio observations by the Event Horizon Telescope  (?). In collisionless magnetorotational turbulence, the half of the energy flux injected via MRI is supposed to be viscously dissipated due to the pressure anisotropy  (?, ?), and the remaining half turns into ion and electron heating at the microscopic scales smaller than the ion Larmor radius. Our previous study of microscopic turbulence using hybrid gyrokinetics showed that, regarding the heating at the microscopic scales, ions are heated more efficiently than electrons when slow-mode-like compressive fluctuations dominate the Alfvénic ones  (?). Thus, the fact that the slow-magnetosonic fluctuations have twice as large energy as the Alfvénic ones indicates that at least half of the energy cascaded down to the ion Larmor scale dissipates into ion heating (see also  (?) for some caveats of this conclusion). Recently, the one dimensional energy transport model of hot accretion disks showed that the preferential ion heating due to the slow-mode-like compressive fluctuations can substantially influence the global temperature distribution  (?). We should note that this discussion assumes that MRI turbulence is active in the accretion flows around M87 and Sgr A*  (?, ?). However, the theoretical examinations of the Event Horizon Telescope data prefer magnetically arrested disk (MAD) regime, in which MRI is supposed to be strongly (but not completely) suppressed due to strong magnetic fields. That being said, the contribution of MRI-driven turbulence to the radiation profile in a MAD state is still being debated. Furthermore, the accumulation and amplification of the poloidal magnetic field due to accretion caused by MRI are crucial for achieving the MAD state  (?, ?). There is also a study that suggests MRI is not even suppressed in MADs, contrary to previous claims  (?).
@@
MRI turbulence can accelerate high-energy non-thermal particles via magnetic reconnection  (?, ?, ?, ?), and these higher energy particles are further accelerated via stochastic acceleration through wave-particle interactions  (?). If non-thermal protons are accelerated to higher energies, these protons can produce cosmic neutrinos seen by IceCube experiments  (?). Previous studies on stochastic acceleration using MHD and test-particle simulations exhibit particle acceleration with a hard-sphere type diffusion coefficient in momentum space  (?, ?, ?). However, these MHD simulations lack sufficient spacial resolution, underestimating the acceleration efficiency at lower energies. Our highest resolution MRI simulation resolves the inertial range of MRI turbulence for the first time. This allows us to evaluate the particle acceleration efficiency at much lower energies down to dissipation scale. This would shed light on modeling particle acceleration inside accretion flows.
@@
Methods
@@
$\frac{\partial\mathbf{u}}{\partial t}+(\mathbf{u}_{0}+\mathbf{u})\cdot\nabla\mathbf{u}=-\nabla P+\mathbf{B}\cdot\nabla\mathbf{B}-2\Omega\hat{\mathbf{z}}\times\mathbf{u}-\mathbf{u}\cdot\nabla\mathbf{u}_{0}$ (1)
@@
where $\mathbf{u}$ is the flow velocity, $\mathbf{B}$ is the magnetic field, $P$ is the total pressure, $\Omega$ is the local angular velocity of the disk, $q=-(\mathrm{d}\ln\Omega/\mathrm{d}\ln r)_{r=r_{0}}$ is the shear rate, and $\mathbf{u}_{0}=-q\Omega x\hat{\mathbf{y}}$ is the background flow. Due to the incompressible condition, the density $\rho$ is spatio-temporally constant. Although MRI turbulence can excite substantial acoustic wave power in accretion flows, the MRI turbulence itself is highly incompressible. In fact, the particle-in-cell simulation demonstrated that the incompressible approximation is valid in collisionless magnetorotational turbulence in a shearing box  (?). We assume that the rotation of the disk is Keplerian, i.e., $q=3/2$. The boundary condition is set to periodic in $y$ and $z$, and to shearing periodic in $x$  (?).
@@
We numerically solve (1)-(3) via pseudo-spectral code Calliope  (?, ?). In order to adopt the pseudospectral method, we enforce the triply periodic boundary conditions by transforming to the shearing coordinate, $y\mapsto y-q\Omega tx$. This transformation makes the radial wavenumber time dependent, $k_{x}(t)=k_{x}+q\Omega tk_{y}$, where $k_{x}$ and $k_{x}(t)$ are the radial wavenumber in the shearing frame and the laboratory frame, respectively. To avoid $k_{x}$ from ever growing, we adopt the remapping method where the fields are mapped to the original non-shearing coordinate every $T=L_{y}/(q\Omega L_{x})$  (?, ?).
@@
The computational domain was discretized into $N_{x}\times N_{y}\times N_{z}=N\times N\times N/2$ grid points. We set $N$ at 256 initially and gradually increased it to 8192 after nonlinear saturation. Each time $N$ was increased, we continued the simulation until the spectral shapes near the dissipation scale did not change before $N$ was increased again. When $N$ was increased to 8192, the simulation was continued for a duration exceeding 100 $\Omega^{-1}$ from the initial time. Thus, the MRI turbulence near the injection scale was sufficiently developed before we start the highest resolution run.
@@
To broaden the inertial range, we used fourth-order hyperviscosity $\nu_{\mathrm{h}}\nabla^{4}\mathbf{u}$ and hyperresistivity $\eta_{\mathrm{h}}\nabla^{4}\mathbf{B}$ to terminate the cascade. Magnetic Prandtl number $\mathrm{Pm}=\nu_{\mathrm{h}}/\eta_{\mathrm{h}}$ was set to unity. We found that when the order of the hyper-dissipation was eight, there appeared an unphysical roll up in the kinetic-energy spectrum.
@@
Acknowledgments
YK thanks Hideaki Miura, Tomohiko Watanabe, and Jim Stone for fruitful comments on this study. Numerical computations reported here were carried out on Fugaku at the RIKEN Center for Computational Science (Project ID: hp220027 and hp230006), on ATERUI II at Center for Computational Astrophysics in National Astronomical Observatory of Japan, on Oakbridge-CX, Oakforest-PACS and Wisteria/BDEC-01 at the University of Tokyo, on Flow at Nagoya University, and on ITO at Kyushu University.
Funding
  • YK is supported by JSPS KAKENHI grant No. 20K14509. SSK is supported by JSPS KAKENHI grant Nos. 22K14028, 21H04487, 23H04899 and the Tohoku Initiative for Fostering Global Researchers for Interdisciplinary Sciences (TI-FRIS) of MEXTs Strategic Professional Development Program for Young Researchers.
Author contributions
  • Conceptualization: YK and SSK
  • Methodology: YK
  • Investigation: YK and SSK
  • Visualization: YK
  • Funding acquisition: YK and SSK
  • Writing – original draft: YK and SSK
  • Writing – review & editing: YK and SSK
Competing interests
  • The authors declare that they have no competing interests.
Data and materials availability
  • All data needed to evaluate the conclusions in the paper are present in the paper and/or the Supplementary Materials.
Supplementary materials
Figs. S1 to S3.
Figure 1: Morphology of finely resolved spatial structures of magnetorotational turbulence. Spatial distributions of (A) flow and (B) magnetic field intensity on $x=0$, $y=0$, and $z=0$ planes, where $x,y$, and $z$ denote radial, azimuthal, and vertical directions, respectively. The white lines are typical magnetic field lines.
Figure 2: Small-scale structure of fluctuations. The same snapshot as Fig. 1, but high-pass filtered at $|\mathbf{k}|\geq 20$ (as shown below, the fluctuations with $|\mathbf{k}|\geq 20$ are in the inertial range). (A) and (B) correspond to the flow and magnetic field intensity distributions, respectively, The bluish pseudo color in (B) indicates the magnitude of unfiltered electric current density $\mathbf{J}$ in a logarithmic scale. The green lines are typical unfiltered magnetic field lines.
Figure 3: Wavenumber anisotropy of magnetic-energy spectrum. Two dimensional contour of magnetic-energy spectrum on (A) $(k_{z},k_{y})$ plane integrated over $k_{x}$ and on (B) $(k_{z},k_{x})$ plane integrated over $k_{y}$. The snapshot was taken immediately after remapping from the shearing coordinate to the laboratory coordinate so that the radial wavenumber $k_{x}$ in the both coordinate coincides. The colorbar applies to both (A) and (B). (C) $k_{z}$-intercept vs. $k_{y}$- and $k_{x}$- intercepts of contour lines in (A) and (B). The gray shaded area indicates to the dissipation range.
Figure 4: Spectra of magnetorotational turbulence. (A) Kinetic, magnetic, and total energy spectra compensated by $k^{3/2}$, where $k$ is the wavenumber. (B) Spectra of $x$, $y$, and $z$ components of magnetic energy. (C) Cross-scale energy flux $\Pi^{f^{<}}_{g^{>}}$ denoting the transfer from the field $f$ (which is the flow field when $f=u$ and is the magnetic field when $f=B$) with the wavenumber smaller than $k$ to the field $g$ with the wavenumber larger than $k$. The gray shaded area indicates to the dissipation range. The arrow indicates the wavenumber $k_{\mathrm{MRI}}=2$ which denotes the wavenumber of the fastest growing MRI modes.
Figure 5: Locality of nonlinear energy transfer. (A)-(D) Shell-to-Shell energy transfer function $\mathcal{T}_{fg}$, representing the transfer from field $f$ (the flow field for $f=u$ and the magnetic field for $f=B$) with wavenumber $Q$ to field $g$ with wavenumber $K$. (E)-(H) Contribution of the energy transfer from the injection range $k\leq 4$.
Supplementary Materials for
Inertial range of magnetorotational turbulence
Yohei Kawazura, Shigeo S. Kimura
*Correspondence to: kawazura@a.utsunomiya-u.ac.jp
This PDF file includes:
  • Figs. S1 to S3
Figure S1: The spatial distribution of the vector components of magnetic and flow fields. (A), (B), and (C) for the magnetic fields $B_{x}$, $B_{y}$, and $B_{z}$ and (D), (E), and (F) for the flow fields $u_{x}$, $u_{y}$, and $u_{z}$. The snapshot was taken at the same time as in Fig. 1. One finds that only $B_{y}$ has intense large-scale structures, which are supposed to be created by the $\Omega$ effect and inverse cascade, as mentioned in the main text.
Figure S2: Time evolution of spectra. The evolution is displayed from $t=0$ to $0.6\Omega^{-1}$ where $t=0$ corresponds to the time when $0.2\Omega^{-1}$ passed after the resolution was increased to $N=8192$.
Figure S3: Comparison between the spectra of slow-magnetosonic fluctuations and parallel fluctuations. (Left) The spectrum of magnetic and kinetic energy of slow-magnetosonic waves $E_{u}^{\mathrm{SW}}=(1+k_{\|}^{2}/k_{\perp}^{2})u_{\|}^{2}/2$ and $E_{B}^{\mathrm{SW}}=(1+k_{\|}^{2}/k_{\perp}^{2})\delta B_{\|}^{2}/2$, respectively, and (Right) the spectrum of $u_{\|}^{2}/2$ and $\delta B_{\|}^{2}/2$ (the same as Fig. 6A). The only difference is at low-$k$ region were the anisotropy $k_{\|}/k_{\perp}\ll 1$ is not yet developed.